• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Roedelbronn

    Log entry time 08:00:44 on February 1,2002

    Entry number 81454

    keyword=Shift summary

    03:45 Rebooted CODA and ran trigsetup. Everything seems to be working OK now.
    Unfortunately, it seems that one of the problems we had was with TS2, the trigger supervisor. There were error messages in the TS2 windows, which was one of the reasons for the reboot. When we tried to replay the run, espace returned errors and could not process the run. Run 1940 seems to be lost for use in online analysis. It may or may not be possible to recover the data offline. Since we don't know, I must assume that run is not usable and take data until we get 4% with the remaining runs. It puts us roughly 30 minutes behind where we'd intended to be.
    05:00 Completed 4A(e) to statistics of ~4%
    05:30 Completed 4A(f)
    05:40 Commenced low current runs. 3.3 uA is a good current for 4A(e*).
    05:50 Beam is down because of a cavity problem. MCC says they should be back in ~10 minutes.
    06:13 Beam is back.
    06:58 Accelerator took beam away for a quantum efficiency measurement.
    07:54 Finished 4A(e*)
    08:00 Switched to LH2 norad for 4A(f*)